# Complexity of Linear Subsequences of $k$-Automatic Sequences

### 07 Mar 2026 v3

@@ 1 Introduction
The $k$-automatic sequences for integers $k\geq 2$ is an interesting class of sequences introduced by Cobham [12, 13] and was studied for more than 50 years; see [2] for a monograph on automatic sequences. A sequence $(h(i))_{i\geq 0}$ is said to be $k$-automatic if there is a deterministic finite automaton with output (DFAO) computing the $i$-th term, given the base-$k$ representation of $i$ as input. A DFAO is a $6$-tuple $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where
@@
For a given automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ there is a unique associated interior sequence ${\bf h}^{\prime}=(h^{\prime}(i))_{i\geq 0}$ with output values from $Q$, defined by taking the minimal DFAO $(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ and replacing $\tau$ by the identity map $\iota$ on $Q$. Here “unique” means up to renaming of the letters. It follows that $(h(i))_{i\geq 0}$ is the image of $(h^{\prime}(i))_{i\geq 0}$ under the coding $q\rightarrow\tau(q)$.
@@
A classical result of Cobham [13, p. 174] states that if $(h(i))_{i\geq 0}$ is $k$-automatic, then so is the linear subsequence $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$. In this paper, we are interested in the state complexity of linear subsequences of automatic sequences and related topics. Our results on the state complexity of linear subsequences expand on the paper by Zantema and Bosma [30] and answer one of their open questions. More details can be found in Section 4.
@@
At first glance, it might seem that results for the state complexity of problems involving msd-first and lsd-first representations would be closely related. While this is true for arithmetic, it is not true for subsequences of automatic sequences. Roughly speaking, this is because a finite-state transducer can carry out addition and multiplication by constants on lsd-first representations, but not msd-first representations.
Boudet and Comon [7] and Wolper and Boigelot [28] also discussed the size of the automaton for an arbitrary formula in Presburger arithmetic. Furthermore, Amat et al. [3] have developed a tool named Patapsco that computes a lower bound for the number of states in the automaton for a given Presburger arithmetic formula. This tool computes a lower bound without actually constructing the automaton.
@@
Remark 1.
In this paper we sometimes provide state complexity and runtime complexity bounds using big-$O$ notation, and sometimes these bounds involve several variables. For these bounds we use the convention that the bounds hold for all sufficiently large values of all the variables.
@@ 2 Background
The state complexity of a formal language is the number of states in the minimal DFA recognizing it and we say the state complexity of an automatic sequence is the number of states in the minimal DFAO generating it.
@@ 3.1 Addition
The automaton for msd-first base $2$ and lsd-first base $2$ input can be created by the methods explained respectively by Wolper and Boigelot [28] and Boudet and Comon [7]. Furthermore, the automaton for msd-first base $2$ can be found, for example, in Hodgson [15] and the automaton for msd-first base $k$ can be found, for example, in Waxweiler [27]; the msd-first base $k$ construction can be easily modified to obtain the lsd-first base $k$ construction.
@@
We also use the following notation: if some condition is described inside brackets, the brackets are Iverson brackets. The Iverson brackets equal $1$ if the condition described inside them is satisfied and equal 0 otherwise. The next theorem appears to be new.
@@
The construction in its simplest form creates an automaton with consecutive double-state levels and the initial single-state level for $c$. The double-state levels are created as a result of computing both the ceiling and floor of an integer divided by $k$ or the floor and ceiling of two consecutive integers divided by $k$ at each step of the construction. The number of times we need to divide $c$ by $k$ and apply floor and ceiling to get to the 0 state corresponds to $|(c)_{k}|$. Therefore, in the simple form, there are $2|(c)_{k}|+1$ states. The automaton $M_{7}$ for base $3$ is an example of the simple form.
However, for some values of $k$ and $c$, some double-state levels in the simple form are changed to single-state levels, causing the number of states to be lower than the simple form. We study these cases, how they are reflected in the theorem’s formula, and why they are the only possible such cases.
@@
If $k$ divides some integer $c^{\prime}$, then $\lfloor c^{\prime}/k\rfloor=\lceil c^{\prime}/k\rceil$. As a result, in this case, instead of the last level being single-state, the last $\nu_{k}(c)+1$ levels are single-state. So Case 1 is covered by the $\nu_{k}(c)$ in the theorem’s formula. The automaton $M_{6}$ in base $3$ is an example of this case.
Case 2: The $(c)_{k}$ starts with $1$, or $k=2$ and $(c)_{2}$ starts with $10$.
@@ 3.2 Multiplication
In addition to providing an msd-first automaton construction for recognizing the relation $x]_{k}+[y]_{k}=[z]_{k}$, Waxweiler [27] provided one for recognizing $n[x]_{k}=[y]_{k}$. Later, Charlier et al. [11] provided a construction for recognizing $n[x]_{k}+c=[y]_{k}$. Furthermore, the automaton for $n[x]_{k}+c=[y]_{k}$ with msd-first base-$2$ input (resp., lsd-first base-$2$ input) can be created by the methods explained by Wolper and Boigelot [28] (resp., Boudet and Comon [7]). The following theorem considers the automaton accepting $n[x]_{k}+c=[y]_{k}$ in the case $c<n$ with both lsd- and msd-first base $k$ input.
@@ 4 Linear Subsequences of Automatic Sequences
Zantema and Bosma [30, Theorem 16] showed that for lsd-first input, if $(h(i))_{i\geq 0}$ is a $k$-automatic sequence generated by a DFAO of $m$ states, then $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $2m$ states. Furthermore, in the same theorem they showed that for msd-first input, the sequence $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $m^{2}$ states and there is an example where this bound is tight. Additionally, they showed [30, Theorem 17] that the same bounds hold for $(h(i-1))_{i\geq 0}$ where $h(0)$ is any chosen letter from the sequence alphabet.
@@
For lsd-first input, for each $m\geq 2$, there exists a $m$-state DFAO $M$ generating the sequence $(h(i))_{i\geq 0}$ such that at least 2m12m-1 states are required in the DFAO $M^{\prime}$ generating $(h(i+1))_{i\geq 0}$.
@@
We use an adaptation of the Myhill-Nerode theorem [17, §3.4] for DFAOs. The adaptation for DFAOs is also used by Zantema and Bosma [30] and the state equivalence for DFAOs is defined by van Spaendonck [25].
@@
    Let $S_{n}(h^{\prime})$ be the set of length-$n$ subwords of the sequence $(h^{\prime}(i))_{i\geq 0}$. We can define $M^{\prime}=(Q^{\prime},\Sigma_{k},\delta^{\prime},q_{0}^{\prime},\Delta,\tau^{\prime})$ as follows:
@@
    Similarly, from Theorem 2, we know there are at most $k(c+1)m^{2}$ states in the automaton.
@@ 4.2 Thue-Morse Sequence
Consider the automaton generating $(t(i+c))_{i\geq 0}$ built by the construction in Theorem 10 . On input $x$ the automaton reaches the state $t(n[x]_{k}),\ldots,t(n[x]_{k}+c)]$ where the output is $t(n[x])$. So the number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ is at most $\rho_{\bf t}(c+1)$.
@@
Let $x,y\in S_{n}$ with $x\not=y$. We need to find a word $z$ such that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$. If $t([x]_{2}+2^{n})\not=t([y]_{2}+2^{n})$, we can choose $z=\lambda$, the empty string.
Otherwise, assume $t([x]_{2}+2^{n})=t([y]_{2}+2^{n})$. Since $x,y$ are distinct, there must be some position $p$ (numbered from left to right, starting with index $1$) at which $x[p]\not=y[p]$. Without loss of generality, assume $x[p]=0$ and $y[p]=1$. Choose $z=0^{p}$. Let $x^{\prime}$ denote the binary representation of $xz]_{2}+2^{n}$ and $y^{\prime}$ denote the binary representation of $yz]_{2}+2^{n}$. Letting $|u|_{1}$ denote the number of $1$’s in a word $u$, we clearly have $|x^{\prime}|_{1}=|x|_{1}+1$, since a new $1$ arises from the addition of $2^{n}$. However, $|y^{\prime}|_{1}=|y|_{1}$, since adding $2^{n}$ causes the $1$ at position $p$ to become 0 and causes the 0 at position p1p-1 to become $1$, while no other bits change (since $y$ has no two consecutive $1$’s). It follows that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$, as desired. Hence all the elements of $S_{n}$ are pairwise inequivalent, and so the minimal DFAO for $(t(i+c))_{i\geq 0}$ has $F_{n+1}$ states.
Finally, from the estimate $F_{n+1}\geq{1\over 2}\alpha^{n}$ (which is easily proved by induction), where $\alpha=(1+\sqrt{5})/2\doteq 1.61803$ is the golden ratio, we see that if $c=2^{n}$, then the minimal DFAO has at least
$F_{n+1}\geq{1\over 2}\alpha^{n}={1\over 2}c^{{\log\alpha}\over{\log 2}}\geq c^{0.694}$
@@ 4.3 Sequences Whose Shifts Have Small State Complexity
We create an automaton $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where $\Sigma=\Delta=\{0,1\}$ with $O(\log c)$ states generating ${\bf p}_{2}$. In order to create the automaton, we need to consider two possibilities for $i$ where $P_{2}(i+c)=1$. Let $c^{\prime}$ be $2^{s}-c$ where $2^{s}$ is the smallest power of $2$ greater than $c$.
@@
Now we provide a construction for $M$. First, define $\delta(q_{0},0)=q_{0}$. Next, we address Case 1. Let $(c^{\prime})=c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$. For $p\in[0,l]$, let $q^{\prime\prime}_{p}$ be a state in the automaton corresponding to $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{p}$ read and for $p\in[0,l-1]$, let $\delta(q^{\prime\prime}_{p},c^{\prime\prime}_{p+1})=q^{\prime\prime}_{p+1}$ and $\delta(q_{0},1)=q^{\prime\prime}_{0}$.
Next, we address Case 2. Let $c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$ be $(c^{\prime})_{2}$ padded with leading 0s so the length is $|(c)_{2}|$. For $p\in[0,|(c)_{2}|-1]$, let $q^{\prime}_{p}$ be a state corresponding to $c^{\prime}_{0}\cdots c^{\prime}_{p}$ read and for $p\in[0,|(c)_{2}|-2]$, let $\delta(q^{\prime}_{p},c^{\prime}_{p+1})=q^{\prime}_{p+1}$. Let $q_{1}$ be the state corresponding to reading the initial $1$s in Case 2. So $\delta(q^{\prime\prime}_{0},1)=q_{1}$, $\delta(q_{1},1)=q_{1}$, and if $c^{\prime}_{0}=1$, then $q_{1}=q^{\prime}_{0}$. Furthermore, we need to add states and transitions so the automaton accepts $1c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$. Let $p^{\prime}$ be the first index starting from 0 where $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$ differs from $1c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$, which can potentially be $l+1$. We set $\delta(q^{\prime\prime}_{p^{\prime}-1},c^{\prime}_{p^{\prime}-1})=q^{\prime}_{p^{\prime}-1}$.
So the DFAO $M=(Q,\Sigma_{2},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows.
@@
The correctness of $M$ is explained above. The number of states in $M$ is $2+(l+1)+(c)_{2}=O(\log c)$. ∎

### 02 Apr 2026 v5

@@ 1 Introduction
The $k$-automatic sequences for integers $k\geq 2$ is an interesting class of sequences introduced by Cobham [12, 13] and has been studied for more than 50 years; see [2] for a monograph on automatic sequences. A sequence $(h(i))_{i\geq 0}$ is said to be $k$-automatic if there is a deterministic finite automaton with output (DFAO) computing the $i$-th term, given the base-$k$ representation of $i$ as input. A DFAO is a $6$-tuple $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where
@@
For any automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ generated by the minimal DFAO $(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$, there is a unique associated interior sequence ${\bf h}^{\prime}=(h^{\prime}(i))_{i\geq 0}$ with output values from the states of the minimal DFAO generating $(h(i))_{i\geq 0}$, defined by taking the minimal DFAO and replacing $\tau$ by the identity map $\iota$ on $Q$. Here “unique” means up to renaming of the letters. Therefore, $(h(i))_{i\geq 0}$ is the image of $(h^{\prime}(i))_{i\geq 0}$ under the coding $q\rightarrow\tau(q)$.
@@
A classical result of Cobham [13, p. 174] states that if $(h(i))_{i\geq 0}$ is $k$-automatic, then so is the linear subsequence $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$. In this paper, we are interested in the state complexity of linear subsequences of automatic sequences and related topics. Our results on the state complexity of linear subsequences expand on the paper by Zantema and Bosma [30] and we answer one of their open questions. More details can be found in Section 4.
@@
At first glance, it might seem that results for the state complexity of problems involving msd-first and lsd-first representations would be closely related. While this is true for arithmetic, it is not true for subsequences of automatic sequences. Roughly speaking, this is because a deterministic finite-state transducer can carry out addition and multiplication by constants on lsd-first representations, but not msd-first representations.
Boudet and Comon [7] and Wolper and Boigelot [28] also discussed the size of the automaton for an arbitrary formula in Presburger arithmetic. Furthermore, Amat et al. [3] have developed a tool named Patapsco that computes a lower bound for the number of states in the automaton for a given Presburger arithmetic formula; this tool computes the lower bound without actually constructing the automaton.
@@ 2 Background
The state complexity of a formal language is defined as the number of states in the minimal DFA recognizing it and we define he state complexity of an automatic sequence as the number of states in the minimal DFAO generating it.
@@ 3.1 Addition
The automaton for msd-first base 22 and lsd-first base $2$ input can be created by the methods explained, respectively, by Wolper and Boigelot [28] and Boudet and Comon [7]. Furthermore, the automaton for msd-first base $2$ can be found, for example, in Hodgson [15] and the automaton for msd-first base $k$ can be found, for example, in Waxweiler [27, Proposition 4.4.2]; the msd-first base $k$ construction can be easily modified to obtain the lsd-first base $k$ construction.
@@
We also use the following notation: if some condition is described inside brackets, the brackets are Iverson brackets. The Iverson bracket $[P]$ equals $1$ if the condition $P$ is satisfied, and 0 otherwise. The next theorem appears to be new.
@@
In its simplest form, our construction creates an automaton with consecutive double-state levels and an initial single-state level for $c$. At each step of the construction, the double-state levels are created by computing both the ceiling and floor of an integer divided by $k$ or the floor and ceiling of two consecutive integers divided by $k$. The number of times we need to divide $c$ by $k$ and apply floor and ceiling to get to the 0 state is $|(c)_{k}|$. Therefore, in the simple form, there are $2|(c)_{k}|+1$ states. The automaton $M_{7}$ for base $3$ is an example of the simple form.
However, for some values of $k$ and $c$, some double-state levels in the simple form collapse to single-state levels, causing the number of states to be lower than the simple form. We study these cases, how they are reflected in Eq. (2), and why they are the only possible such cases.
@@
If $k$ divides some integer $c^{\prime}$, then $\lfloor c^{\prime}/k\rfloor=\lceil c^{\prime}/k\rceil$. As a result, in this case, instead of the last level being single-state, the last $\nu_{k}(c)+1$ levels are single-state. So Case 1 is covered by the $\nu_{k}(c)$ in Eq. (2). The automaton $M_{6}$ in base $3$ is an example of this case.
Case 2: The word $(c)_{k}$ starts with $1$, or $k=2$ and $(c)_{2}$ starts with $10$.
@@ 3.2 Multiplication
In addition to providing an msd-first automaton construction for recognizing the relation $x]_{k}+[y]_{k}=[z]_{k}$, Waxweiler [27, Proposition 4.4.3] provided one for recognizing $n[x]_{k}=[y]_{k}$. Later, Charlier et al. [11] provided a construction for recognizing $n[x]_{k}+c=[y]_{k}$. Furthermore, the automaton for $n[x]_{k}+c=[y]_{k}$ with msd-first base-$2$ input (resp., lsd-first base-$2$ input) can be created by the methods explained by Wolper and Boigelot [28] (resp., Boudet and Comon [7]). The following theorem considers the automaton accepting $n[x]_{k}+c=[y]_{k}$ in the case $c<n$ with both lsd- and msd-first base $k$ input.
@@ 4 Linear Subsequences of Automatic Sequences
Zantema and Bosma [30, Theorem 16] showed that for lsd-first input, if $(h(i))_{i\geq 0}$ is a $k$-automatic sequence generated by a DFAO of $m$ states, then $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $2m$ states. Furthermore, in the same theorem they showed that for msd-first input, the sequence $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $m^{2}$ states and there is an example where this bound is tight. Additionally, they [30, Theorem 17] showed that the same bounds hold for $(h(i-1))_{i\geq 0}$ where $h(0)$ is any chosen letter from the sequence alphabet.
@@
For lsd-first input, for each $m\geq 2$, there exists an $m$-state DFAO $M$ generating the sequence $(h(i))_{i\geq 0}$ such that at least 2m12m-1 states are required in the DFAO $M^{\prime}$ generating $(h(i+1))_{i\geq 0}$.
@@
We use an adaptation of the Myhill-Nerode theorem [16, §3.4] for DFAOs. The adaptation for DFAOs is also used by Zantema and Bosma [30] and the state equivalence for DFAOs is defined by van Spaendonck [25].
@@
    Let $S_{n}(h^{\prime})$ be the set of length-$n$ subwords of the sequence $(h^{\prime}(i))_{i\geq 0}$. We define $M^{\prime}=(Q^{\prime},\Sigma_{k},\delta^{\prime},q_{0}^{\prime},\Delta,\tau^{\prime})$ as follows:
@@
    Similarly, the states are length-$(c+1)$ subwords of the interior sequence $(h^{\prime}(i))_{i\geq 0}$.
@@ 4.2 Thue-Morse Sequence
Consider the automaton generating $(t(i+c))_{i\geq 0}$ built by the construction in Theorem 9 . On input $x$ the automaton reaches the state $t([x]_{k}),\ldots,t([x]_{k}+c)]$ where the output is $t([x]+c)$. So the number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ is at most $\rho_{\bf t}(c+1)$.
@@
Let $x,y\in S_{n}$ with $x\not=y$. We need to find a word $z$ such that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$. If $t([x]_{2}+2^{n})\not=t([y]_{2}+2^{n})$, we can choose $z=\lambda$, the empty word.
Otherwise, assume $t([x]_{2}+2^{n})=t([y]_{2}+2^{n})$. Since $x,y$ are distinct, there must be some position $p$ (numbered from left to right, starting with index $1$) at which $x[p]\not=y[p]$ where $w[i]$ is the $i$-th letter of $w$. Without loss of generality, assume $x[p]=0$ and $y[p]=1$. Choose $z=0^{p}$. Let $x^{\prime}$ denote the binary representation of $xz]_{2}+2^{n}$ and $y^{\prime}$ denote the binary representation of $yz]_{2}+2^{n}$. Letting $|u|_{1}$ denote the number of $1$’s in a word $u$, we clearly have $|x^{\prime}|_{1}=|x|_{1}+1$, since a new $1$ arises from the addition of $2^{n}$. However, $|y^{\prime}|_{1}=|y|_{1}$, since adding $2^{n}$ causes the $1$ at position $p$ to become 0 and causes the 0 at position p1p-1 to become $1$, while no other bits change (since $y$ has no two consecutive $1$’s). It follows that $t([xz]_{2}+2^{n})\not=t([yz]_{2}+2^{n})$, as desired. Hence all the elements of $S_{n}$ are pairwise inequivalent, and so the minimal DFAO for $(t(i+c))_{i\geq 0}$ has $F_{n+1}$ states.
Finally, from the estimate $F_{n+1}\geq\frac{1}{2}\alpha^{n}$ (which is easily proved by induction), where $\alpha=(1+\sqrt{5})/2\doteq 1.61803$ is the golden ratio, we see that if $c=2^{n}$, then the minimal DFAO has at least
$F_{n+1}\geq\frac{1}{2}\alpha^{n}=\frac{1}{2}c^{\frac{\log\alpha}{\log 2}}\geq c^{0.694}$
@@ 4.3 Sequences Whose Shifts Have Small State Complexity
We create an automaton $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where $\Sigma=\Delta=\{0,1\}$ with $O(\log c)$ states generating $(P_{2}(i+c))_{i\geq 0}$. In order to create the automaton, we need to consider two possibilities for $i$ where $P_{2}(i+c)=1$. Let $c^{\prime}$ be $2^{s}-c$ where $2^{s}$ is the smallest power of $2$ greater than $c$. For now let us ignore the case where $c$ is a power of $2$.
@@
Now we provide a construction for $M$. Let $\delta(q_{0},0)=q_{0}$. We first address Case 1. Let $(c^{\prime})=c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$. For $p\in[0,l]$, let $q^{\prime\prime}_{p}$ be a state in the automaton corresponding to $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{p}$ read and for $p\in[0,l-1]$, let $\delta(q^{\prime\prime}_{p},c^{\prime\prime}_{p+1})=q^{\prime\prime}_{p+1}$ and $\delta(q_{0},1)=q^{\prime\prime}_{0}$. Furthermore, let $\tau(q^{\prime\prime}_{l})=1$.
Next, we address Case 2. Let $c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$ be $(c^{\prime})_{2}$ padded with leading 0s so the length is $|(c)_{2}|$. For $p\in[0,|(c)_{2}|-1]$, let $q^{\prime}_{p}$ be a state corresponding to $c^{\prime}_{0}\cdots c^{\prime}_{p}$ read and for $p\in[0,|(c)_{2}|-2]$, let $\delta(q^{\prime}_{p},c^{\prime}_{p+1})=q^{\prime}_{p+1}$. Let $q_{1}$ be the state corresponding to reading the initial $1$s in Case 2. So $\delta(q_{1},1)=q_{1}$ and $\delta(q_{1},c^{\prime}_{0})=q^{\prime}_{0}$. We know $c^{\prime}_{0}=0$ if $c$ is not a power of $2$. Furthermore, let $\tau(q^{\prime}_{|(c)_{2}|-1})=1$.
Now we need to define transitions so the automaton is deterministic and accepts $1^{j}c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$ for all $j\geq 1$. If there are no 0s in $c_{0}^{\prime\prime}\cdots c^{\prime\prime}_{l}$, let $p^{\prime\prime}=l$; otherwise, let $p^{\prime\prime}$ be the first index starting from 0 where $c^{\prime\prime}_{p^{\prime\prime}+1}=0$. We set $\delta(q^{\prime\prime}_{p^{\prime\prime}},1)=q_{1}$; this handles $j\geq p^{\prime\prime}+2$. For $p\in[0,p^{\prime\prime}-1]$, we set $\delta(q^{\prime\prime}_{p},0)=q^{\prime}_{0}$; this handles $1\leq j^{\prime}\leq p^{\prime\prime}$. Let $p^{\prime}$ be the first index starting from 0 where $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$ differs from $1^{p^{\prime\prime}+1}c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$, which can potentially be $l+1$. We set $\delta(q^{\prime\prime}_{p^{\prime}-1},c^{\prime}_{p^{\prime}-p^{\prime\prime}-1})=q^{\prime}_{p^{\prime}-p^{\prime\prime}-1}$; this handles $j=p^{\prime\prime}+1$.
So if $c$ is not a power of 22, the DFAO $M=(Q,\Sigma_{2},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows.
@@
The correctness of $M$ is explained above. The number of states in $M$ is $2+(l+1)+(c)_{2}=O(\log c)$.
If $c$ is a power of $2$, then $c=c^{\prime}$ and it is easy to see that the same idea works with less complications because $|(c)_{2}|=|(c^{\prime})_{2}|$. We just need to modify the automaton defined above and set $q^{\prime\prime}_{i}=q^{\prime}_{i}$ for $i\in[0,l]$, $q_{1}=q^{\prime\prime}_{0}$, $\tau(q_{0})=1$ and omit the transitions involving $p^{\prime}$ and $p^{\prime\prime}$. The number of states will remain $O(\log c)$. ∎